OpenStudy (anonymous):

A bag contains 9 green marbles and 11 white marbles. You select a marble at random. What are the odds in favor of picking a green marble? a.9:20 b.2:9 c.11:9 d.9:11

OpenStudy (anonymous):

a

OpenStudy (anonymous):

prob(green) = 9/20 prob(not green) = 11/20 odds in favour of green = (9/20) : (11/20) = 9: 11

OpenStudy (anonymous):

I think It Is D

OpenStudy (anonymous):

thanks @Dubba2000 and @Kazuma

OpenStudy (anonymous):

Np :D

OpenStudy (anonymous):

i say a because there are 20 marble and 9 green so it would be A

OpenStudy (anonymous):

and welcome lol